how many liters of baking soda must be added to 240 liters of a 20% baking soda solution to get a 60% baking soda solution
Answer by Maths68(1474)

Soluton A
Amount = x
Concentration = 100%=1
================================================================================
Solution B
Amount = 240 liters
Concentration = 20%=0.2
================================================================================
Final Solution
Amount=240+x
Concentration=60%=0.6
================================================================================
(Amount of A * Concentration of A)+(Amount of B * Concentration of B) = (Final Amount * Final Concentration)
(x*1)+(240*0.2)=(240+x)*0.6
x+48=144+0.6x
x-0.6x=144-48
0.4x/0.4=96/0.4
x=240
================================================================================
240 liters of baking soda must be added to 240 liters of a 20% baking soda solution to get a 60% baking soda solution